# FlashKAT: Understanding and Addressing Performance Bottlenecks in the Kolmogorov-Arnold Transformer

### v2

@@ 4 FlashKAT
In the previous section, we identified that the KAT has a memory bottleneck in the backward pass of the group-wise rational function, a discovery that previous approaches had not recognized (yang2024kolmogorov; molina2019pad; delfosse2021recurrent). In this section, we first outline the computations required for this operation, then provide an analysis of the standard group-wise rational function used in the state-of-the-art KAT to pin down the source of the bottleneck, and finally present FlashKAT, our proposed Kolmogorov-Arnold Transformer that builds on an optimized group-wise rational function that efficiently leverages gradient accumulation and memory hierarchy to minimize global memory accesses.
@@
$\frac{\partial F}{\partial b_{j}}=-x^{j}\frac{A(x)}{|A(x)|}\frac{P(x)}{Q(x)^{2}},$ (8)
$\frac{\partial F}{\partial x}=\frac{\partial P(x)}{\partial x}\frac{1}{Q(x)}-\frac{\partial Q(x)}{\partial x}\frac{P(x)}{Q^{2}(x)}.$ (9)
Equations 7, 8, and 9 provides the element-wise gradients; however, since each group of size $d_{g}$, each batch of size $B$ and each element in the sequence of size $N$ share an identical set of coefficients, the contributions of each must be accumulated on $\frac{\partial F}{\partial a_{i}}$ and $\frac{\partial F}{\partial b_{j}}$. Let us denote $\frac{\partial F}{\partial a_{g,i}}$ and $\frac{\partial F}{\partial b_{g,j}}$ as the gradient for coefficient $i$ and $j$ of group $g$. Then we can express this accumulation in Equations 10 and 11:
$\frac{\partial F}{\partial a_{g,i}}=\sum^{B}_{i^{\prime}=1}\sum^{N}_{j^{\prime}=1}\sum^{d_{g}}_{k^{\prime}=1}\frac{\partial F(X_{i^{\prime},j^{\prime},g\cdot d_{g}+k^{\prime}})}{\partial a_{g,i}},$ (10)
$\frac{\partial F}{\partial b_{g,j}}=\sum^{B}_{i^{\prime}=1}\sum^{N}_{j^{\prime}=1}\sum^{d_{g}}_{k^{\prime}=1}\frac{\partial F(X_{i^{\prime},j^{\prime},g\cdot d_{g}+k^{\prime}})}{\partial b_{g.j}}.$ (11)
@@
Provided an input $\mathbf{X}\in\mathbb{R}^{B\times N\times d}$, an output gradient $\mathbf{dO}\in\mathbb{R}^{B\times N\times d}$ and coefficients $\mathbf{A}\in\mathbb{R}^{n_{g}\times m}$ and $\mathbf{B}\in\mathbb{R}^{n_{g}\times n}$, we want to compute the element-wise gradients $\frac{\partial F}{\partial\mathbf{A}}$ ($\mathbf{dA}$), $\frac{\partial F}{\partial\mathbf{B}}$ ($\mathbf{dB})$, and $\frac{\partial F}{\partial\mathbf{X}}$ ($\mathbf{dX}$). In Algorithm 1, we provide the standard method for computing each of these gradients in the KAT. The algorithm is structured sequentially, and all the on-chip computation is done in the subroutine calls, following Equations 7, 8, and 9, for readability.
@@
4:Initialize $\textbf{dA}\leftarrow 0\in\mathbb{R}^{n_{g}\times m}$, $\mathbf{dB}\leftarrow 0\in\mathbb{R}^{n_{g}\times n}$, and $\textbf{dX}\leftarrow 0\in\mathbb{R}^{B\times N\times d}$.
@@
6:  Load $\mathbf{X}_{i}$ from global memory
7:  Load $\mathbf{dO}_{i}$ from global memory
8:  for $1\leq j\leq S_{\text{block}}$ do
9:  $k\leftarrow\lfloor((i\cdot T+j)\mod d)/d_{g}\rfloor$
10:  Load $\mathbf{A}_{k,1:m}$ and $\mathbf{B}_{k,1:n}$ from global memory
11:  $\mathbf{dX}_{i,j}\leftarrow\textbf{Comp\_dX}(\mathbf{A}_{k,1:m},\mathbf{B}_{k,1:n},\mathbf{X}_{i,j},\mathbf{dO}_{i,j})$
12:  $\mathbf{dA}^{\prime}_{k,1:m}\leftarrow\textbf{Comp\_dA}(\mathbf{A}_{k,1:m},\mathbf{B}_{k,1:n},\mathbf{X}_{i,j},\mathbf{dO}_{i,j})$
13:  $\mathbf{dB}^{\prime}_{k,1:n}\leftarrow\textbf{Comp\_dB}(\mathbf{A}_{k,1:m},\mathbf{B}_{k,1:n},\mathbf{X}_{i,j},\mathbf{dO}_{i,j})$
@@
14:  Load $\mathbf{dA}_{k,1:m}$ from global memory, set $\mathbf{dA}_{k,1:m}\leftarrow\mathbf{dA}_{k,1:m}+\mathbf{dA}^{\prime}_{k,1:m}$, write back to global memory
@@
15:  Load $\mathbf{dB}_{k,1:n}$ from global memory, set $\mathbf{dB}_{k}\leftarrow\mathbf{dB}_{k}+\mathbf{dB}^{\prime}_{k}$, write back to global memory
16:  end for
17:  Write back $\mathbf{dX}_{i}$ to global memory
18:end for
19:return $\mathbf{dA},\mathbf{dB},\mathbf{dX}$
@@
At a high level, Algorithm 1 iterates over the $T$ blocks in the grid, where each block has $S_{\text{block}}$ threads that compute the associated gradients for $\mathbf{dA}$, $\mathbf{dB}$, $\mathbf{dX}$. In the case of $\mathbf{dA}$ and $\mathbf{dB}$, for each of these computations, the thread performs an atomic add, which requires reading from global memory, accumulating the value, and then storing the result back to global memory as an indivisible unit of work. However, since each of these threads must perform an atomic add for each of the $m+n$ coefficients, the atomic adds must occur sequentially. As such, for each of these atomic adds, a stall occurs. Since multiple threads within the same block and across different blocks may be writing to the same location, there will be extreme resource contention. Such resource contention inevitably leads to high latency and stalls without utilizing the available bandwidth, as we demonstrated in Section 3.
@@
Let us analyze the total number of global memory accesses incurred by Algorithm 1. We can see that each $\mathbf{X}_{i}$ and $\mathbf{dO}_{i}$ requires $S_{\text{block}}$ reads from global memory. Then writing $\mathbf{dX}_{i}$ back to global memory requires $S_{\text{block}}$ writes to global memory. Since these 3 memory actions occur $T$ times they require $3TS_{\text{block}}=3\frac{B\cdot N\cdot d}{S_{\text{block}}}\cdot S_{\text{block}}=3\cdot B\cdot N\cdot d$ global memory accesses. Regarding the coefficients, each element in the output $dX$ requires loading $m+n$ coefficients from global memory. It also requires the aforementioned atomic add, which includes an additional $(m+n)$ reads and $(m+n)$ writes to global memory. Thus, since there are $T$ blocks of $S_{\text{block}}$ threads, there are $TS_{\text{block}}\cdot 3(m+n)=3(m+n)\cdot B\cdot N\cdot d$ accesses. As a result, there are $3(m+n+1)\cdot B\cdot N\cdot d$ global memory accesses in total. To reiterate, the memory accesses for the coefficients for the atomic add are much worse due to the additional contention with other threads.
@@
We now present an optimized method for computing the backward pass of the group-wise rational function in Algorithm 2. To reduce global memory accesses, we restructure the grid from 1D to 2D. Such a restructuring mirrors the format of Equation 10, allowing an easier mapping of it to the algorithm. The first dimension of our reconstructed grid is allocated for resolving the $B$ and $N$ summations, whereas the second dimension is responsible for the $d_{g}$ summation. The size of this first dimension in the grid is $B\cdot N/S_{\text{block}}$, whereas the size of the second dimension is $n_{g}$ (the total number of groups in GR-KAN).
@@
4:Divide $\mathbf{A}$ and $\mathbf{B}$ into $n_{g}$ blocks $\mathbf{A_{j}}\in\mathbb{R}^{m}$ and $\mathbf{B}_{j}\in\mathbb{R}^{n}$.
5:Initialize $\mathbf{dA}\leftarrow 0\in\mathbb{R}^{n_{g}\times m}$, $\mathbf{dB}\leftarrow 0\in\mathbb{R}^{n_{g}\times n}$, and $\mathbf{dX}\leftarrow 0\in\mathbb{R}^{B\times N\times d}$.
@@
8:  Load $\mathbf{A}_{j}$, $\mathbf{B_{j}}$ from global memory
9:  Load $\mathbf{X}_{i,j}$ from global memory
10:  Load $\mathbf{dO}_{i,j}$ from global memory
11:  Initialize $\mathbf{dA_{j}}^{\prime}\leftarrow 0$ and $\mathbf{dB_{j}}^{\prime}\leftarrow 0$
12:  for $1\leq k\leq S_{\text{block}}$ do
13:   for $1\leq l\leq d_{g}$ do
14:   $\mathbf{dX}_{i,j,k,l}\leftarrow\textbf{Comp\_dX}(\mathbf{A}_{j,1:m},\mathbf{B}_{j,1:n},\mathbf{X}_{i,j,k,l},\mathbf{dO}_{i,j,k,l})$
15:   $\mathbf{dA}^{\prime}_{j}\leftarrow\mathbf{dA}^{\prime}_{j}+\textbf{Comp\_dA}(\mathbf{A}_{j,1:m},\mathbf{B}_{j,1:n},\mathbf{X}_{i,k},\mathbf{dO}_{i,j,k,l})$
16:   $\mathbf{dB}^{\prime}_{j}\leftarrow\mathbf{dB}^{\prime}_{j}+\textbf{Comp\_dB}(\mathbf{A}_{j,1:m},\mathbf{B}_{j,1:n},\mathbf{X}_{i,j,k,l},\mathbf{dO}_{i,j,k,l})$
17:   end for
18:  end for
19:  Load $\mathbf{dA}_{j}$ from global memory, set $\mathbf{dA}_{j}\leftarrow\mathbf{dA}_{j}+\mathbf{dA}^{\prime}_{j}$, write back to global memory
20:  Load $\mathbf{dB}_{j}$ from global memory, set $\mathbf{dB}_{j}\leftarrow\mathbf{dB}_{j}+\mathbf{dB}^{\prime}_{j}$, write back to global memory
21:  Write back $\mathbf{dX}_{i,j}$ to global memory
22:  end for
23:end for
24:return $\mathbf{dA},\mathbf{dB},\mathbf{dX}$
@@
We now analyze the total global memory accesses for Algorithm 2. Each block first loads one set of coefficients from the global memory associated with the group to which the block is assigned. Since there is only 1 set of coefficients per block, this action only requires $m+n$ accesses. Then the algorithm loads the associated $\mathbf{X}_{i,j}$ and $\mathbf{dO}_{i,j}$ from global memory, each of which is of size $S_{\text{block}}\cdot d_{g}$. After computing $\mathbf{dX}_{i,j}$, its $S_{\text{block}}\cdot d_{g}$ values are written back to global memory. Then, an atomic add loads $m+n$ values from $\mathbf{dA}_{j}$ and $\mathbf{dB_{j}}$, accumulates them with $\mathbf{dA}^{\prime}_{j}$ and $\mathbf{dB}^{\prime}_{j}$, and writes them back to global memory. Since all these operations occur across the $T\cdot n_{g}$ blocks, there are $T\cdot n_{g}(3S_{\text{block}}\cdot d_{g}+3(m+n))=3\frac{B\cdot N}{S_{\text{block}}}\cdot\frac{d}{d_{g}}(S_{\text{block}}\cdot d_{g}+(m+n))=3(\frac{m+n}{S_{\text{block}}\cdot d_{g}}+1)\cdot B\cdot N\cdot d$ global memory access. Compared to Algorithm 1, we can see that Algorithm 2 reduces the number of global memory accesses and subsequently the number of atomic adds by a factor of $\frac{1}{S_{\text{block}}\cdot d_{g}}$.
@@ 5 Results
Using our experimental configurations from Section 3, we verify the efficacy of Algorithm 2 compared to Algorithm 1. We implement Algorithm 2 in Triton (10.1145/3315508.3329973) and run our experiments on a 4060 Ti. First, we compare the backward pass of the FlashKAT group-wise rational function against the KAT group-wise rational function by measuring cycles, time, and throughput utilization for SMs and memory. As shown in Table 3, the FlashKAT group-wise rational function achieves a speedup of 140.5x over the KAT group-wise rational function, with improvements across the utilization of SM, cache, and memory, contributing to the reduction in backward pass time.
@@
Figure 3: The warp‐state statistics for the FlashKAT group-wise rational function backward pass. ”Computing - Selected” is where the warp does useful computation; all the others are memory stall (MS) states.
Secondly, we again conduct the warp state statistics analysis for the backward pass of the FlashKAT group-wise rational function. From observing the warp state statistics in Figure 3, we can see that only the ”Stall Long Scoreboard” is still longer than the ”Selected” compute state; all the other 8 memory stalls are shorter than compute. Furthermore, the ”Stall Long Scoreboard” has reduced from 981.51 cycles to 2.31 cycles. Such a result indicates that the new kernel is now spending more time performing computation rather than mostly waiting for memory transfers.
@@
We validate that the training capabilities of FlashKAT are consistent with KAT by training it on ImageNet-1k for image recognition (russakovsky2015imagenet). While training, we also measure the training throughput of FlashKAT to compare against KAT, ViT, and DeiT. The data loader time is excluded from this measurement as it is dependent on external factors (e.g., CPU, main memory, etc.). We measure the throughput in images/second as the average of 100 samples, and provide the 95% confidence intervals.
Our training procedure follows an identical curriculum to (yang2024kolmogorov), wherein we follow the hyperparameters of DeiT (touvron2021training). As such, we train using a batch size of 1024 with the AdamW optimizer (loshchilov2017decoupled). For the models, we apply Mimetic initialization to the attention layers (trockman2023mimetic) and set the patch size to 16. In GR-KAN, each of the 8 groups shares its 6 numerator coefficients and has 4 unique denominator coefficients. The first layer of GR-KAN has its group-wise rational function initialized to the identity function, and the second layer is initialized to a Swish function (ramachandran2017searching). For data augmentation and regularization techniques, we apply RandAugment (cubuk2020randaugment), Mixup (zhang2017mixup), CutMix (yun2019cutmix), Random Erasing (zhong2020random), weight decay, Label Smoothing (szegedy2016rethinking) and Stochastic Depth (huang2016deep). We conduct all experiments using an H200 GPU. The Appendix provides the complete set of model configurations and training hyperparameters.
@@
In addition to a significant speedup, the proposed FlashKAT also offers reduced rounding errors when computing the gradients for the group-wise rational function coefficients. This is a byproduct of our method, accumulating gradients in each block through a reduction sum rather than atomic adds. Table 5 demonstrates that the mean absolute error in gradient outputs in FlashKAT is orders of magnitude smaller than KAT. As such, FlashKAT is less susceptible to rounding errors that may contribute toward unstable gradient updates (e.g. low-precision training). Due to space limitations, we provide more details in the Appendix.

### v3

@@ 4 FlashKAT
In the previous section, we identified that KAT has a memory bottleneck in the backward pass of the group-wise rational function, a finding that prior approaches had not recognized (Yang and Wang 2024; Molina et al. 2019; Delfosse et al. 2020, 2021). In this section, we first outline the computations required for this operation, then provide an analysis of the standard group-wise rational function used in KAT to pin down the source of the bottleneck, and finally present FlashKAT, our proposed Kolmogorov-Arnold Transformer that builds on an optimized group-wise rational function that efficiently leverages gradient accumulation and memory hierarchy to minimize global memory accesses.
@@
$\frac{\partial F}{\partial b_{j}}=-x^{j}\frac{A(x)}{|A(x)|}\cdot\frac{P(x)}{Q(x)^{2}},$ (8)
$\frac{\partial F}{\partial x}=\frac{\partial P(x)}{\partial x}\cdot\frac{1}{Q(x)}-\frac{\partial Q(x)}{\partial x}\cdot\frac{P(x)}{Q^{2}(x)}.$ (9)
Equations 7, 8, and 9 provides the local element-wise gradients; however, since each batch of size $B$, each element in the sequence of size $N$, and each group of size $d_{g}$ share an identical set of coefficients, the contributions of each must be accumulated on $\frac{\partial F}{\partial a_{i}}$ and $\frac{\partial F}{\partial b_{j}}$. Let us denote $\frac{\partial F}{\partial a_{g,i}}$ and $\frac{\partial F}{\partial b_{g,j}}$ as the gradient for coefficient $i$ and $j$ of group $g$ and $\frac{\partial\mathcal{L}}{\partial F}$ the upstream gradient. Then we can express this accumulation in Equations 10 and 11:
$\frac{\partial\mathcal{L}}{\partial a_{g,i}}=\sum^{B}_{i^{\prime}=1}\sum^{N}_{j^{\prime}=1}\sum^{d_{g}}_{k^{\prime}=1}\frac{\partial\mathcal{L}}{\partial F^{i^{\prime},j^{\prime}}_{g\cdot d_{g}+k^{\prime}}}\cdot\frac{\partial F^{i^{\prime},j^{\prime}}_{g\cdot d_{g}+k^{\prime}}}{\partial a_{g,i}},$ (10)
$\frac{\partial\mathcal{L}}{\partial b_{g,j}}=\sum^{B}_{i^{\prime}=1}\sum^{N}_{j^{\prime}=1}\sum^{d_{g}}_{k^{\prime}=1}\frac{\partial\mathcal{L}}{\partial F^{i^{\prime},j^{\prime}}_{g\cdot d_{g}+k^{\prime}}}\cdot\frac{\partial F^{i^{\prime},j^{\prime}}_{g\cdot d_{g}+k^{\prime}}}{\partial b_{g,j}}.$ (11)
@@
Provided an input $\mathbf{X}\in\mathbb{R}^{B\times N\times d}$, an upstream gradient $\mathbf{dO}\in\mathbb{R}^{B\times N\times d}$ and coefficients $\mathbf{A}\in\mathbb{R}^{n_{g}\times m+1}$ and $\mathbf{B}\in\mathbb{R}^{n_{g}\times n}$, we want to compute the gradients $\mathbf{dA}\in\mathbb{R}^{n_{g}\times m+1}$ ($\frac{\partial\mathcal{L}}{\partial\mathbf{A}}$), $\mathbf{dB}\in\mathbb{R}^{n_{g}\times n}$ ($\frac{\partial\mathcal{L}}{\partial\mathbf{B}}$), and $\mathbf{dX}\in\mathbb{R}^{B\times N\times d}$ ($\frac{\partial\mathcal{L}}{\partial\mathbf{X}}$). In Algorithm 1, we provide the method for computing each of these gradients in KAT. The algorithm is structured sequentially, and all on-chip computation is performed within subroutine calls, with the upstream gradients applied to Equations 7, 8, and 9 for readability.
@@
4:Initialize $\textbf{dA}\leftarrow 0\in\mathbb{R}^{n_{g}\times m+1}$, $\mathbf{dB}\leftarrow 0\in\mathbb{R}^{n_{g}\times n}$, and $\textbf{dX}\leftarrow 0\in\mathbb{R}^{B\times N\times d}$.
@@
6:  Load $\mathbf{X}_{i}$ and $\mathbf{dO}_{i}$ from global memory
7:  for $1\leq j\leq S_{\text{block}}$ do
8:  $k\leftarrow\lfloor(((i-1)\cdot S_{\text{block}}+j)\mod d)/d_{g}\rfloor$
9:  Load $\mathbf{A}_{k,:}$ and $\mathbf{B}_{k,:}$ from global memory
10:  $\mathbf{dX}_{i,j}\leftarrow\textbf{Comp\_dX}(\mathbf{A}_{k,:},\mathbf{B}_{k,:},\mathbf{X}_{i,j},\mathbf{dO}_{i,j})$
11:  $\mathbf{dA}^{\prime}_{k,:}\leftarrow\textbf{Comp\_dA}(\mathbf{A}_{k,:},\mathbf{B}_{k,:},\mathbf{X}_{i,j},\mathbf{dO}_{i,j})$
12:  $\mathbf{dB}^{\prime}_{k,:}\leftarrow\textbf{Comp\_dB}(\mathbf{A}_{k,:},\mathbf{B}_{k,:},\mathbf{X}_{i,j},\mathbf{dO}_{i,j})$
@@
13:  Load $\mathbf{dA}_{k,:}$ from global memory, $\mathbf{dA}_{k,:}\leftarrow\mathbf{dA}_{k,:}+\mathbf{dA}^{\prime}_{k,:}$, write back $\mathbf{dA}_{k,:}$ to global memory $\triangleright$ atomic add
@@
14:  Load $\mathbf{dB}_{k,:}$ from global memory, $\mathbf{dB}_{k,:}\leftarrow\mathbf{dB}_{k,:}+\mathbf{dB}^{\prime}_{k,:}$, write back $\mathbf{dB}_{k,:}$ to global memory $\triangleright$ atomic add
15:  end for
16:  Write back $\mathbf{dX}_{i}$ to global memory
17:end for
18:return $\mathbf{dA},\mathbf{dB},\mathbf{dX}$
@@
At a high level, Algorithm 1 iterates over the $T$ blocks in the grid, where each block has $S_{\text{block}}$ threads that compute the associated gradients for $\mathbf{dA}$, $\mathbf{dB}$, $\mathbf{dX}$. In the case of $\mathbf{dA}$ and $\mathbf{dB}$, for each of these computations, the thread performs an atomic add, which requires reading from global memory, accumulating the value, and then storing the result back to global memory as an indivisible unit of work. However, since each of these threads must perform an atomic add for each of the $m+n+1$ coefficients, the atomic adds must occur sequentially. As such, for each of these atomic adds, a stall occurs. Since multiple threads within the same block and across different blocks may write to the same location, there will be severe resource contention. Such resource contention inevitably leads to high latency and stalls without utilizing the available bandwidth, as we demonstrated in Section 3.
@@
Let us analyze the total number of global memory accesses incurred by Algorithm 1. We can see that each $\mathbf{X}_{i}$ and $\mathbf{dO}_{i}$ requires $S_{\text{block}}$ reads from global memory. Then writing $\mathbf{dX}_{i}$ back to global memory requires $S_{\text{block}}$ writes to global memory. Since these 3 memory actions occur $T$ times they require $3TS_{\text{block}}=3\frac{B\cdot N\cdot d}{S_{\text{block}}}\cdot S_{\text{block}}=3\cdot B\cdot N\cdot d$ global memory accesses. Regarding the coefficients, each element in the output $dX$ requires loading $m+n+1$ coefficients from global memory. It also requires the aforementioned atomic add, which includes an additional $m+n+1$ reads and writes to global memory. Thus, since there are $T$ blocks of $S_{\text{block}}$ threads, there are $TS_{\text{block}}\cdot 3(m+n+1)=3(m+n+1)\cdot B\cdot N\cdot d$ accesses. As a result, there are $3(m+n+2)\cdot B\cdot N\cdot d$ global memory accesses in total. To reiterate, memory accesses for atomic add operations for coefficient accumulations are much worse due to additional contention from other threads.
@@
We now present an optimized method for computing the backward pass of the group-wise rational function in Algorithm 2. To reduce global memory accesses, we restructure the grid from 1D to 2D. Such a restructuring mirrors the format of Equation 10, facilitating a more straightforward mapping of it to the algorithm. The first dimension of our reconstructed grid is allocated for resolving the $B$ and $N$ summations, whereas the second dimension is responsible for the $d_{g}$ summation. The size of this first dimension in the grid is $B\cdot N/S_{\text{block}}$, whereas the size of the second dimension is $n_{g}$ (the total number of groups in GR-KAN).
@@
4:Divide $\mathbf{A}$ and $\mathbf{B}$ into $n_{g}$ blocks $\mathbf{A}_{j}\in\mathbb{R}^{m+1}$ and $\mathbf{B}_{j}\in\mathbb{R}^{n}$.
5:Initialize $\mathbf{dA}\leftarrow 0\in\mathbb{R}^{n_{g}\times m+1}$, $\mathbf{dB}\leftarrow 0\in\mathbb{R}^{n_{g}\times n}$, and $\mathbf{dX}\leftarrow 0\in\mathbb{R}^{B\times N\times d}$.
@@
8:  Load $\mathbf{A}_{j}$, $\mathbf{B}_{j}$, $\mathbf{X}_{i,j}$, and $\mathbf{dO}_{i,j}$ from global memory
9:  Initialize $\mathbf{dA}_{j}^{\prime}\leftarrow 0$ and $\mathbf{dB}_{j}^{\prime}\leftarrow 0$
10:  for $1\leq k\leq S_{\text{block}}$ do
11:   for $1\leq l\leq d_{g}$ do
12:  $\mathbf{dX}_{i,j,k,l}\leftarrow\textbf{Comp\_dX}(\mathbf{A}_{j},\mathbf{B}_{j},\mathbf{X}_{i,j,k,l},\mathbf{dO}_{i,j,k,l})$
13:  $\mathbf{dA}^{\prime}_{j}\leftarrow\mathbf{dA}^{\prime}_{j}+\textbf{Comp\_dA}(\mathbf{A}_{j},\mathbf{B}_{j},\mathbf{X}_{i,j,k,l},\mathbf{dO}_{i,j,k,l})$
14:  $\mathbf{dB}^{\prime}_{j}\leftarrow\mathbf{dB}^{\prime}_{j}+\textbf{Comp\_dB}(\mathbf{A}_{j},\mathbf{B}_{j},\mathbf{X}_{i,j,k,l},\mathbf{dO}_{i,j,k,l})$
15:   end for
16:  end for
17:  Load $\mathbf{dA}_{j}$ from global memory, $\mathbf{dA}_{j}\leftarrow\mathbf{dA}_{j}+\mathbf{dA}^{\prime}_{j}$, write back $\mathbf{dA}_{j}$ to global memory $\triangleright$ atomic add
18:  Load $\mathbf{dB}_{j}$ from global memory, $\mathbf{dB}_{j}\leftarrow\mathbf{dB}_{j}+\mathbf{dB}^{\prime}_{j}$, write back $\mathbf{dB}_{j}$ to global memory $\triangleright$ atomic add
19:  Write back $\mathbf{dX}_{i,j}$ to global memory
20:   end for
21:end for
22:return $\mathbf{dA},\mathbf{dB},\mathbf{dX}$
@@
We now analyze the total global memory accesses for Algorithm 2. Each block first loads one set of coefficients from the global memory associated with the group to which the block is assigned. Since there is only 1 set of coefficients per block, this operation requires only $m+n+1$ accesses. Then the algorithm loads the associated $\mathbf{X}_{i,j}$ and $\mathbf{dO}_{i,j}$ from global memory, each of size $S_{\text{block}}\cdot d_{g}$. After computing $\mathbf{dX}_{i,j}$, its $S_{\text{block}}\cdot d_{g}$ values are written back to global memory. Then, an atomic add loads $m+n+1$ values from $\mathbf{dA}_{j}$ and $\mathbf{dB_{j}}$, accumulates them with $\mathbf{dA}^{\prime}_{j}$ and $\mathbf{dB}^{\prime}_{j}$, and writes them back to global memory. Since all these operations occur across the $T\cdot n_{g}$ blocks, there are $T\cdot n_{g}(3S_{\text{block}}\cdot d_{g}+3(m+n+1))=3(\frac{m+n+1}{S_{\text{block}}\cdot d_{g}}+1)\cdot B\cdot N\cdot d$ global memory access. Compared to Algorithm 1, we can see that Algorithm 2 reduces the number of global memory accesses and atomic adds by a factor of $\frac{1}{S_{\text{block}}\cdot d_{g}}$.
@@ 5 Results
Using our experimental configurations from Section 3, we verify the efficacy of Algorithm 2 compared to Algorithm 1. We implement Algorithm 2 in Triton (Tillet et al. 2019) and run our experiments on a 4060 Ti. First, we compare the backward pass of the FlashKAT group-wise rational function with that of the KAT group-wise rational function by measuring cycles, time, and throughput utilization for SMs and memory. As shown in Table 3, the FlashKAT group-wise rational function achieves a speedup of 140.5x over the KAT group-wise rational function, with improvements across SM, cache, and memory utilization, reducing backward pass time.
@@
Figure 3: The warp‐state statistics for the FlashKAT group-wise rational function backward pass. “Computing - Selected” is where the warp does useful computation; all the others are memory stall (MS) states.
Secondly, we again conduct the warp state statistics analysis on the backward pass of the FlashKAT group-wise rational function. From observing the warp state statistics in Figure 3, we can see that only the “Stall Long Scoreboard” is still longer than the “Selected” compute state; all the other 8 memory stalls are shorter than compute. Furthermore, the “Stall Long Scoreboard” has reduced from 981.51 cycles to 2.31 cycles. This result indicates the new kernel spends more time performing computations than waiting for memory transfers.
@@
We validate that FlashKAT’s training capabilities are consistent with KAT by training it on ImageNet-1k (Russakovsky et al. 2015). During training, we measure FlashKAT’s training throughput to compare it with KAT, ViT, and DeiT. The data loader time is excluded from this measurement because it depends on external factors (e.g., CPU, main memory, etc.). We measure throughput in images/second as the average of 100 samples, and provide 95% confidence intervals.
Our training procedure follows the same curriculum as (Yang and Wang 2024), using the hyperparameters of DeiT (Touvron et al. 2021). As such, we train using a batch size of 1024 with the AdamW optimizer (Loshchilov and Hutter 2017). For the models, we apply Mimetic initialization to the attention layers (Trockman and Kolter 2023) and set the patch size to 16. In GR-KAN, each of the 8 groups shares its 6 numerator coefficients and has 4 unique denominator coefficients. The first layer of GR-KAN has its group-wise rational function initialized to the identity function, and the second layer is initialized to a Swish function (Ramachandran et al. 2017). For data augmentation and regularization techniques, we apply RandAugment (Cubuk et al. 2020), Mixup (Zhang et al. 2017), CutMix (Yun et al. 2019), Random Erasing (Zhong et al. 2020), weight decay, Label Smoothing (Szegedy et al. 2016) and Stochastic Depth (Huang et al. 2016). We conduct all experiments using an H200 GPU. The Appendix provides the complete set of model configurations and training hyperparameters.
@@
In addition to a significant speedup, the proposed FlashKAT also reduces rounding errors when computing gradients for the group-wise rational function coefficients. This is a byproduct of our method, accumulating gradients in each block through a reduction sum rather than atomic adds. Table 5 demonstrates that the mean absolute error in gradient outputs in FlashKAT is orders of magnitude smaller than in KAT. As such, FlashKAT is less susceptible to rounding errors that may contribute toward unstable gradient updates (e.g., low-precision training). Due to space limitations, we provide more details in the Appendix.